# Strongly Interacting Dark Matter admixed Neutron Stars

### v1

@@ 3.2 Stability
Not every solution of the TOV equations yield a stable neutron star. Small perturbations in the metric field should settle back to the original solution. By solving a Sturm-Liouville problem one can identify the stable solutions by demanding the following criterion Hippert:2022snq ; Pitz:2024xvh ; Barbat:2024yvi
@@
where $N$ is the number of particles of the two types of fluids in the star. This problem can be translated to demanding that both Eigenvalues of the equation in (3) are positive. The number of particles can be calculated simultaneously to the TOV equations via
$\frac{dN}{dr}=4\pi\left(1-\frac{2m}{r}\right)^{-1/2}nr^{2}dr.$ (4)
@@ 3.3 Tidal Deformability
When two compact objects orbit each other, they get deformed by the gravitational tidal field of one another. The tidal deformability $\Lambda$ is a quantity that describes the proportionality between the tidal field and the induced quadrupole deformation. We will work with the unit-less version of the tidal deformability that is connected to the dimensionful tidal deformability $\lambda$ and the dimensionless second Love number $k_{2}$ via
@@
In the late stages of an inspiral event the tidal deformability induces a gravitational wave phase given by a combination of the tidal deformabilities of the two participating neutron stars. To leading order in the tidal deformability this phase can be determined by the parameter $\tilde{\Lambda}$ as LIGO_GW170817
@@
where $\Lambda_{1}$ and $\Lambda_{2}$ are the tidal deformabilities and $M_{1}$ and $M_{2}$ are the total masses of the two neutron stars. As this cannot separate both $\Lambda_{i}$, we choose conservatively the extreme cases that either one entirely dominates for the limits in $\Lambda$ below.
@@ 4 Results
Figure 2: The investigated set of central pressures of ordinary matter ($p_{0,O}$) and dark matter ($p_{0,D}$). Gray points indicate stable solutions outside the constraints chosen in (12). Colored points indicate stable solutions that meet the constraints at $>$1% dark matter fraction for the color coding and $<$1% in yellow. The dark matter mass increases from 0.37 GeV (top) and 1 GeV (middle) to 4 GeV (bottom). The left-hand side shows the results for ordinary matter equation of state I, and the right hand side for III. Both are for the light dark matter equation of state.
For this intermediate mass range, we solved the two-fluid TOV equation by iterating over both the central pressures of ordinary matter and dark matter and perform the stability analysis. In figure 2 we show the range of central pressures that we investigated. Gray points indicate a stable solution to the TOV-equation. One immediate observation is that the addition of dark matter makes it possible to sustain larger central pressures for ordinary matter. This is particularly interesting as this grants access to new regimes of equations of state that could potentially have interesting phenomenology. Moreover, heavier dark matter results in larger possible central pressures which can also be inferred from scaling arguments (see section 3.1).
As an additional constraint, we want to investigate under which circumstances the objects considered in this paper are compatible with astronomical observation. Because this is an exploratory study, we intentionally choose loose bounds on the mass, the radius and the tidal deformability given by
@@
Figure 5: The observable radius of the ordinary $R_{O}$ in comparison to the radius of the dark matter radius $R_{D}$ in km in the same color-coding as figure 2. The left-hand side shows the results for ordinary matter equation of state I, and the right hand side for III. Both are for the light dark matter equation of state. The dotted line indicates where $R_{O}=R_{D}$. Solutions above indicate a dark matter halo scenario while solutions below indicate a dark matter core scenario.
We show contributions from the individual parts to the mass in figure 4 and for the radius in figure 5. We again also illustrate if the dark matter contribution to the total mass is limited to 1% of the total mass. The first important observation from figures 3-5 is that changing the dark matter mass by an order of magnitude has much more severe impact than changing the equation of state, except for the radius. The observable (maximal) radius is basically entirely dominated by the ordinary matter equation of state.
@@ 5 Summary
In this work we studied the impact of a strongly-interacting dark matter component on neutron star properties, using G2-QCD. The equation of state for the dark fluid was obtained at finite density from first principles using lattice field theory Maas:2012wr ; Wellegehausen:2013cya ; Hajizadeh:2017jsw . We used the two equations of state at our disposal, differing in their mass scales, and found that they have similar influence on the resulting properties of the mixed star. On the standard model side, we covered a large range of possibilities for the ordinary matter equation of state coming from the interpolation between low and high densities using piecewise polytropes Kurkela:2014vha .
As is well known from investigations of the ordinary matter equations of state, they all result in different mass-radius-relations. The addition of dark matter has similar effects on all three of them. For each combination of equations of state we varied the mass of the dark matter candidate and the central pressure of the ordinary fluid and dark fluid and performed a stability analysis. The addition of dark matter, in general, allows for a larger possible central pressures which should be kept in mind when investigating equations of state. As expected, we find that light dark matter has a stronger effect on the properties of the neutron star and that by tuning the ratio of central pressures one is also able to find solutions that fulfill the experimental bounds. For some combination of the dark matter candidate mass and the ratio of central pressures we find results that do not resemble neutron stars ($M_{DM}$/$M_{tot}$>1%) but could provide an explanation for non-standard compact stellar objects. An astrophysical observation of exotic compact objects also in gravitational wave experiments would allow for the discrimination of the composition of such objects. We leave this for future explorations.
@@ 6.1 Results for the heavy dark matter equation of state
In this appendix, we compile a set of results for the heavy equation of state for the dark matter sector, as well as for the equation of state II. We did not use them in the main text as they do not substantially change the overall picture. This can be seen for the mass-radius relation shown in figures 6 and 7, which need to be compared with figure 3.

### v2

@@ 3.2 Stability
Not every solution of the TOV equations yields a stable neutron star. Small perturbations in the metric field should settle back to the original solution. By solving a Sturm-Liouville problem one can identify the stable solutions by demanding the following criterion Hippert:2022snq ; Pitz:2024xvh ; Barbat:2024yvi
@@
where $N_{O/D}$ is the total number of ordinary/dark particles and $\varepsilon_{c,O/D}$ is the energy density of ordinary/dark matter in the center of the star. This problem can be translated into demanding that both the eigenvalues of the equation in (3) be positive. The number of particles can be calculated simultaneously to the TOV equations via
$dN=4\pi\left(1-\frac{2m}{r}\right)^{-1/2}nr^{2}dr.$ (4)
@@ 3.3 Tidal Deformability
When two compact objects orbit each other, they get deformed by the gravitational tidal field of one another. The tidal deformability $\Lambda$ is a quantity that describes the proportionality between the tidal field and the induced quadrupole deformation. We will work with the unitless version of the tidal deformability that is connected to the dimensionful tidal deformability $\lambda$ and the dimensionless second Love number $k_{2}$ via
@@
In contrast to black holes, neutron stars have a non-vanishing tidal deformability. As a result, the gravitational wave signal of a neutron star binary differs from the one produced by a black hole merger by a (potentially time-dependent) phase in the late stages of the inspiral. To leading order in the tidal deformabilities of the two neutron stars, this phase can be determined by the parameter $\tilde{\Lambda}$ as LIGO_GW170817
@@
where $\Lambda_{1}$ and $\Lambda_{2}$ are the tidal deformabilities and $M_{1}$ and $M_{2}$ are the total masses of the two neutron stars. Thus, deviations in this quantity compared to an ordinary matter neutron star give a measure of the impact a dark matter component will have on the gravitational wave signal of two compact stellar objects. As the observed phase difference cannot separate both $\Lambda_{i}$, we choose conservatively the extreme cases that either one entirely dominates for the limits in $\Lambda$ below.
@@ 4 Results
Figure 2: The investigated set of central pressures of ordinary matter ($p_{0,O}$) and dark matter ($p_{0,D}$). Gray points indicate stable solutions outside the constraints chosen in (12). Colored points indicate stable solutions that meet the constraints at >>1% dark matter fraction for the color coding and <<1% in yellow. The dark matter mass increases from 0.37 GeV (top) and 1 GeV (middle) to 4 GeV (bottom). The left-hand side shows the results for ordinary matter equation of state I, and the right hand side for III. Both are for the light dark matter equation of state.
For this intermediate mass range, we solved the two-fluid TOV equation by iterating over both the central pressures of ordinary matter and dark matter and perform the stability analysis. In figure 2 we show the range of central pressures that we investigated. Gray points indicate a stable solution to the TOV-equation. One immediate observation is that the addition of dark matter makes it possible to sustain larger central pressures for ordinary matter as was also found in Kain:2021hpk ; Valdez-Alvarado:2012rct . This is particularly interesting as this grants access to new regimes of equations of state that could potentially have interesting phenomenology. Moreover, heavier dark matter results in larger possible central pressures which can also be inferred from scaling arguments (see section 3.1).
As an additional constraint, we want to investigate under which circumstances the objects considered in this paper are compatible with astronomical observation. Given the recent developments, e. g. the observation of neutron star candidates with possibly surprising features Salmi:2024bss ; Doroshenko:2022nwp , we intentionally choose very relaxed bounds, to accommodate possible future discoveries. This will also allow our exploratory study to remain valid during possible future such observations. This leads to the following choice of bounds:
@@
Figure 5: The observable radius of the ordinary matter $R_{O}$ in comparison to the radius of the dark matter $R_{D}$ in km in the same color-coding as figure 2. The left-hand side shows the results for ordinary matter equation of state I, and the right hand side for III. Both are for the light dark matter equation of state. The dotted line indicates where $R_{O}=R_{D}$. Solutions above indicate a dark matter halo scenario while solutions below indicate a dark matter core scenario.
We show contributions from the individual parts to the mass in figure 4 and for the radius in figure 5. We again also illustrate if the dark matter contribution to the total mass is limited to 1% of the total mass. The first important observation from figures 3-5 is that changing the dark matter mass by an order of magnitude has much more severe impact than changing the equation of state, except for the radius. The observable radius is basically entirely dominated by the ordinary matter equation of state.
@@ 5 Summary
In this work we studied the impact of a strongly-interacting dark matter component on neutron star properties, using G2-QCD. The equation of state for the dark fluid was obtained at finite density from first principles using lattice field theory Maas:2012wr ; Wellegehausen:2013cya ; Hajizadeh:2017jsw . We used the two equations of state at our disposal, differing in their mass scales, and found that they have similar influence on the resulting properties of the mixed star. On the standard model side, we covered a range of equations of state that capture the current constraints on the mass, radius and tidal deformability. They were used in Kurkela:2014vha by interpolating between low and high densities using piecewise polytropes.
As is well known from investigations of the ordinary matter equations of state, they all result in different mass-radius-relations. The added dark matter has similar effects on all three of them. For each combination of equations of state we varied the mass of the dark matter candidate and the central pressure of the ordinary fluid and dark fluid and performed a stability analysis. The addition of dark matter, in general, allows for a larger possible central pressures which should be kept in mind when investigating equations of state Kain:2021hpk ; Valdez-Alvarado:2012rct . As expected, we find that light dark matter has a stronger effect on the properties of the neutron star and that by tuning the ratio of central pressures one is also able to find solutions that fulfill the experimental bounds. For some combination of the dark matter candidate mass and the ratio of central pressures we find results that do not resemble neutron stars ($M_{DM}$$M_{tot}$>1%) but could provide an explanation for non-standard compact stellar objects. An astrophysical observation of exotic compact objects also in gravitational wave experiments would motivate a more careful analysis of these objects. We leave this for future exploration.
@@ 6 Appendix
6.1 Method of interpolation using polytropes
Figure 6: Labelling of the parameters of the piecewise polytropes
In order to use the lattice results for a calculation of the TOV equation we have to convert them to pressure $p$ and energy density $\varepsilon$. In order to ensure thermodynamic consistency at the level of guaranteeing $n\mu=p\varepsilon$, we use the framework of piecewise polytropes to interpolate between the data points
Following Read:2008iy , the polytropic equation of state is defined by Kurkela:2014vha
$p=Kn^{\Gamma}$ (13)
where $n$ is the number density and $\Gamma$ is called the adiabatic index. The energy density is fixed by
$\begin{split}d\left(\frac{\varepsilon}{n}\right)&=-p\,\,d\left(\frac{1}{n}\right)\\ \varepsilon&=cn+\frac{K}{(\Gamma-1)}n^{\Gamma},\end{split}$ (14)
where in the second line an integration has been performed, introducing an integration constant $c$. From $n\mu=p+\varepsilon$ follows for the chemical potential.
$\mu=c+\frac{\Gamma}{\Gamma+1}Kn^{\Gamma-1}$ (15)
This formalism is used to interpolate between data points, see figure 6 for an illustration. The $i$’th piecewise polytrope, valid from $n_{i-1}$ to $n_{i}$, depends on the parameters $c_{i}$, $K_{i}$ and $\Gamma_{i}$. In the following, we will denote the functions depending on the number density and the $i$’th parameters as $p_{i}(n)$, $\varepsilon_{i}(n)$ and $\mu_{i}(n)$, and the values of these functions at $n_{i}$ as $p_{i}$, $\varepsilon_{i}$, $\mu_{i}$.
At low-density follows from $\lim_{n\to 0}\mu=m_{c}$ that for the first polytrope, $c_{0}=m_{c}$. The other assumption is that at very low density $\Gamma_{0}=\frac{5}{3}$ holds, which describes a free Fermi gas. This parameterization coincides with $n=c(\mu^{2}-m_{c}^{2})^{\frac{3}{2}}$ at low densities. Demanding continuity, one can use the following formulas to find the parameters of the next piece of the polytrope,
$\begin{split}\mu_{i+1}-\mu_{i}&=\mu_{i+1}(n_{i+1})-\mu_{1}(n_{i})\\ &=c_{i+1}+K_{i+1}\frac{\Gamma_{i+1}}{\Gamma_{i+1}-1}n_{i+1}^{\Gamma_{i+1}-1}-c_{i+1}-K_{i+1}\frac{\Gamma_{i+1}}{\Gamma_{i+1}-1}n_{i}^{\Gamma_{i+1}-1}\\ &=K_{i+1}\frac{\Gamma_{i+1}}{\Gamma_{i+1}-1}(n_{i+1}^{\Gamma_{i+1}-1}-n_{i}^{\Gamma_{i+1}-1})\\ &=K_{i}n_{i}^{\Gamma_{i}-\Gamma_{i+1}}\frac{\Gamma_{i+1}}{\Gamma_{i+1}-1}(n_{i+1}^{\Gamma_{i+1}-1}-n_{i}^{\Gamma_{i+1}-1})\end{split}$ (16)
where in the last line we have eliminated the yet unknown $K_{i+1}$ by demanding continuity in the pressure $p_{i}(n_{i})=p_{i+1}(n_{i})$. The following expression can be solved by a zero search algorithm to find $\Gamma_{i+1}$
$\Gamma_{i+1}K_{i}n_{i}^{\Gamma_{i}-\Gamma_{i+1}}=\frac{\mu_{i+1}-\mu_{i}}{n_{i+1}^{\Gamma_{i+1}}-n_{i}^{\Gamma_{i+1}}}(\Gamma_{i+1}-1).$ (17)
$K_{i+1}$ and $c_{i+1}$ are then obtained from continuity in pressure and chemical potential, respectively. We will also need the pressure at the start of the next polytrope for the next iteration, which is given by the polytropic equation.
$\begin{split}K_{i+1}&=n_{i}^{-\Gamma_{i+1}}P_{i}\\ c_{i+1}&=\mu_{i+1}-\frac{K_{i+1}n_{i+1}^{\Gamma_{i+1}-1}\Gamma_{i+1}}{\Gamma_{i+1}-1}\\ P_{i+1}&=K_{i+1}n_{i+1}^{\Gamma_{i+1}}.\end{split}$ (18)
For the first polytrope, $K_{0}$ has to be calculated once from the first data point of $\mu$ and $n$ via
$K_{0}=(\mu_{0}-c_{0})\frac{\Gamma_{0}-1}{\Gamma_{0}}n_{0}^{1-\Gamma_{0}}$ (19)
Finally, the speed of sound is given by
$c_{s,i}^{2}=\frac{dp}{d\varepsilon}=\frac{dp}{dn}\left(\frac{d\varepsilon}{dn}\right)^{-1}=\frac{\Gamma_{i}p}{p+\varepsilon}.$ (20)
As is visible, this implies usually that the speed of sound is discontinuous, when switching from one polytrope to the next.
In principle, this could again be cured, by combining more data points to fix the derivatives, introducing new parameters. This can be escalated, however this will always yield a discontinuity at latest at the $n$’th derivative for $n$ data points. This could only be avoided by a global fit, using a continuous function. This would, however, also eliminate phase transitions. Eventually, there is no optimal choice with having only a finite number of data points. Since the TOV equations do not depend on the speed of sound, and do not require it to be continuous, but only the (guaranteed) continuity and compatibility of the equation of state, polytropes are the minimal solution to guarantee this feature. The slight kinks are integrated over, and thus an over estimation and under estimation by the kinked solutions will average, to some degree, out. Hence, we choose this option here, following Kurkela:2014vha ; Dengler:2021qcq ; Pitz:2024xvh . The stability analysis in section 3.2 uses only integrated quantities and initial conditions, and is therefore not sensitive to local kinks in the equation of state.
We apply this framework for both the ordinary matter equations of state and the dark matter equation of state.
@@ 6.2 Results for the heavy dark matter equation of state
In this appendix, we compile a set of results for the heavy equation of state for the dark matter sector, as well as for the equation of state II. We did not use them in the main text as they do not substantially change the overall picture. This can be seen for the mass-radius relation shown in figures 7 and 8, which need to be compared with figure 3.